NvidiaIntely (nvidiaintely):

The relationship between volume and temperature of a gas, when pressure and moles of a gas are held constant, is: V*T = k. True False

jhonyy9 (jhonyy9):

k is a constant - yes ?

jhonyy9 (jhonyy9):

bc. than i remember it right V = kT

jhonyy9 (jhonyy9):

in case when the pressure and moles of a gas are held constant

OpenStudy (sweetburger):

Volume and temperature are directly proportional

NvidiaIntely (nvidiaintely):

So its false then?

jhonyy9 (jhonyy9):

yes is false bc. ho i ve wrote above than i remember it right from my study V = kT so from this result T = V/k and k= V/T so the above wrote relation of V*T = k is false sure

NvidiaIntely (nvidiaintely):

Thanks!